# Towards Practical Single-shot Motion Synthesis

### v1

@@ Abstract
Despite the recent advances in the so-called “cold start” generation from text prompts, their needs in data and computing resources, as well as the ambiguities around intellectual property and privacy concerns pose certain counterarguments for their utility. An interesting and relatively unexplored alternative has been the introduction of unconditional synthesis from a single sample, which has led to interesting generative applications. In this paper we focus on single-shot motion generation and more specifically on accelerating the training time of a Generative Adversarial Network (GAN). In particular, we tackle the challenge of GAN’s equilibrium collapse when using mini-batch training by carefully annealing the weights of the loss functions that prevent mode collapse. Additionally, we perform statistical analysis in the generator and discriminator models to identify correlations between training stages and enable transfer learning. Our improved GAN achieves competitive quality and diversity on the Mixamo benchmark when compared to the original GAN architecture and a single-shot diffusion model, while being up to ×$\times 6.8$ faster in training time from the former and ×$\times 1.75$ from the latter. Finally, we demonstrate the ability of our improved GAN to mix and compose motion with a single forward pass. Code available at https://moverseai.github.io/single-shot
@@ 2 Prior Art
Single-shot 3D content generation. More recently, scarce efforts have been made to demonstrate single sample generation to 3D content. Sin3DM [43] trains a diffusion model on a single sample, and leverages and intermediate latent representation to overcome memory and runtime performance issues. SinGRAF [39] bridges a neural rendering representation to accomplish single sample variation generation of specific 3D scenes. For 3D motion single sample generation there exist two approaches, GANimator [28] and SinMDM [33], using adversarial training and denoising diffusion respectively. GANimator consists of 7 stages of skeleton-aware convolutions [1] forming 4 pyramid levels (2-2-2-1 stages) emulating the pyramidal design of SinGAN in the temporal domain, where each stage learns to generate a sequence of different length (up to the input’s one). Changing the core of the aforementioned works, Raab et al. [33] present a motion diffusion model that learns to generate variations of a single motion sequence. SinMDM follows the structure of the UNet-based diffusion model presented in [30] but with a significant detail; they add shift-invariant local attention layers [2] to decrease the receptive field of UNet and avoid overfitting in the single sample training scenario. GANimator is slower to train but significantly faster when generating samples, whereas SinMDM exploits mini-batch training to reduce training time but requires an iterative diffusion process at inference. Further, SinMDM was shown to support more applications without requiring retraining, an important advantage from a practical point of view.
@@ 3.1 Background
where $\mathbb{P}$ denotes a learned distribution and $\tilde{T}_{i}=\alpha\hat{T}_{i}+(1-\alpha)T_{i}$ is a linear combination of the generated and ground truth motion features, respectively. Note that the gradient penalty regularization in Eq. 3.1 enforces Lipschitz continuity and stabilizes the training. To prevent mode collapse due to the single-sample training the $G(\cdot)$ are additionally supervised by an $L1$ reconstruction loss:
@@
where $\lambda_{adv},\lambda_{rec},\lambda_{con}$ as well as $\lambda_{GP}$ from Eq. 3.1 are responsible for weighting the contribution of each loss to the adversarial game of single-sample motion learning. Li et al. [28] discuss the contribution of each loss to the final result, while our focus is concentrated in the balancing between the losses.
@@ 3.3 Cross-Stage Transfer Learning
In addition to increasing convergence via larger mini-batch training, the hierarchical structure of single sample GANs can be exploited to improve convergence rates. ConSinGAN [15] trained multiple stages in parallel, and re-used the discriminator from the previous training stage to improve performance by transferring its weights to initialize the next stage discriminator. GANimator [28] already trains two stages in the same pyramid level but does not perform discriminator transfer across stages or pyramid levels, even though the same discriminator is used – in terms of architecture and capacity, similar to ConSinGAN and in contrast to SinGAN that increases discriminator capacity at each stage.
@@
Our results, visualized in Fig. 4, show that despite the hierarchical approach using the same model and capacity for each stage $S$, most layers lie at low similarity scores. Yet we find that the early layer representations across the corresponding stages of each pyramid level, exhibit higher levels of similarity. Contrary to ConSinGAN that outputs features at each level apart from the last, GANimator reconstructs the output motion at each level. Therefore the early convolutional layers operate on similar features and, as indicated by their CKA scores, extract similar representations, whereas the latter layers apply the level’s motion motifs, style and details.
Based on this analysis, we design a generator transfer learning scheme across levels and stages where each generator’s ($G^{i}_{j}$) early layer weights are initialized from the early layer weights of the corresponding generator of the previous level ($G^{i-1}_{j-2}$).
This ensures that each next training stage is closer to the converged state, thus allowing us to reduce the number of iterations significantly, further boosting single sample GAN training time.

### v2

@@ Abstract
Despite the recent advances in the so-called “cold start” generation from text prompts, their needs in data and computing resources, as well as the ambiguities around intellectual property and privacy concerns pose certain counterarguments for their utility. An interesting and relatively unexplored alternative has been the introduction of unconditional synthesis from a single sample, which has led to interesting generative applications. In this paper we focus on single-shot motion generation and more specifically on accelerating the training time of a Generative Adversarial Network (GAN). In particular, we tackle the challenge of GAN’s equilibrium collapse when using mini-batch training by carefully annealing the weights of the loss functions that prevent mode collapse. Additionally, we perform statistical analysis in the generator and discriminator models to identify correlations between training stages and enable transfer learning. Our improved GAN achieves competitive quality and diversity on the Mixamo benchmark when compared to the original GAN architecture and a single-shot diffusion model, while being up to ×$\times 6.8$ faster in training time from the former and ×$\times 1.75$ from the latter. Finally, we demonstrate the ability of our improved GAN to mix and compose motion with a single forward pass. Project page available at https://moverseai.github.io/single-shot
@@ 2 Prior Art
Single-shot 3D content generation. More recently, scarce efforts have been made to demonstrate single sample generation to 3D content. Sin3DM [43] trains a diffusion model on a single sample, and leverages and intermediate latent representation to overcome memory and runtime performance issues. SinGRAF [39] bridges a neural rendering representation to accomplish single sample variation generation of specific 3D scenes. For 3D motion single sample generation there exist two approaches, GANimator [28] and SinMDM [33], using adversarial training and denoising diffusion respectively. GANimator consists of 7 stages of skeleton-aware convolutions [1] forming 4 pyramid levels (2-2-2-1 stages) emulating the pyramidal design of SinGAN in the temporal domain, where each stage learns to generate a sequence of different length (up to the input’s one). Changing the core of the aforementioned works, Raab et al. [33] present a motion diffusion model that learns to generate variations of a single motion sequence. SinMDM follows the structure of the UNet-based diffusion model presented in [30] but with a significant detail; they add shift-invariant local attention layers [2] to decrease the receptive field of UNet and avoid overfitting in the single sample scenario. GANimator is slower to train but significantly faster when generating samples, whereas SinMDM exploits mini-batch training to reduce training time but requires an iterative diffusion process at inference. Further, SinMDM supports more applications without requiring retraining, an important advantage from a practical point of view.
@@ 3.1 Background
where $\mathbb{P}$ denotes a learned distribution and $\tilde{T}_{i}=\alpha\hat{T}_{i}+(1-\alpha)T_{i}$ is a linear combination of the generated and ground truth motion features, respectively. The last term of the equation, i.e. the gradient penalty regularization, enforces Lipschitz continuity and stabilizes the training. To prevent mode collapse due to the single-sample training the $G(\cdot)$ are additionally supervised by an $L1$ reconstruction loss:
@@
where $\lambda_{adv},\lambda_{rec},\lambda_{con}$, and $\lambda_{GP}$ are responsible for weighting the contribution of each loss to the adversarial game of the single sample learning. Li et al. [28] discuss the contribution of each loss to the final result, while our focus is concentrated in the balancing between the losses.
@@ 3.3 Cross-Stage Transfer Learning
Apart from increasing convergence through larger mini-batch training, the hierarchical structure of single sample GANs can be exploited to improve convergence. ConSinGAN [15] trained multiple stages in parallel and re-used the discriminator from the previous stage to improve performance by transferring its weights to initialize the next stage discriminator. GANimator [28] already trains two stages in the same pyramid level but does not perform discriminator transfer across stages or pyramid levels, even though the same discriminator is used – in terms of architecture and capacity - as in ConSinGAN and in contrast to SinGAN that increases discriminator capacity at each stage.
@@
Our results, depicted in Fig. 4, show that despite the hierarchical approach of using the same model and capacity for each stage, most layers exhibit low similarity scores. Yet we find that the early layer representations across the stages of each pyramid level, exhibit higher levels of similarity. Contrary to ConSinGAN that outputs features at each level apart from the last, GANimator reconstructs the output motion at each level. Therefore the early convolutional layers operate on similar features and, as indicated by their CKA scores, extract similar representations, whereas the latter layers apply the level’s motion motifs, style and details. Based on this analysis, we design a generator transfer learning scheme across levels and stages where each generator’s ($G^{i}_{j}$) early layer weights are initialized from the early layer weights of the previous level generator ($G^{i-1}_{j-2}$). This ensures that each next training stage is closer to the converged state allowing us to reduce the number of iterations significantly, further boosting training time.
@@ 5 Conclusion
Acknowledgements. This work was supported by EU’s Horizon Europe Programme project EMIL-XR [GA 101070533].